# Open-world Machine Learning: A Review and New Outlooks

### v2

@@ 5 Class-incremental Learning
Key Challenges. The central challenge of CIL is the training data of old classes is assumed to be unavailable due to memory limits or privacy issues [119, 120] when learning new classes. However, without accessing old training data, directly updating the model on new classes would cause the catastrophic forgetting problem [121, 14], i.e., the classification accuracy on old classes drops dramatically after learning new classes. By taking a closer look at catastrophic forgetting in CIL, we find that there are mainly three issues, and each of them represents an aspect of forgetting, while the concurrence of them results in the performance degradation of old classes after learning new classes (as shown in Figure 5 (A)): (1) Representation and classifier mismatch. Directly updating the model on new classes could dramatically change the learned representations and class weights for old classes, which leads to the mismatch between old classes’ features and weights. (2) Representation confusion. After learning new classes, the feature extractor would no longer be suitable for old classes. As a result, the input from old and new classes would be confused and overlapped in the deep feature space. (3) Classifier imbalance. The class weights of old and new classes can not be jointly optimized because of the lack of old data. Therefore, the class weights of old and new classes are imbalanced, and the decision boundary can easily shift toward novel classes. The above analysis would not only give a deeper understanding of catastrophic forgetting in CIL but also provide a unified perspective to understand existing methods. Despite the grand challenges, considerable progress has been made to enable the continual ability of DNNs recently. As illustrated in Figure 5 (B), existing methods can be grouped into four main categories: regularization based, exemplar replay, feature replay based and prompt-based methods. In what follows, we review the principles and limitations of different approaches.
5.1 Regularization based CIL
@@
Figure 5: Illustrations of (A) the problem setting and (B) Key challenges of class-incremental learning. Existing technique methods mainly include exemplar replay, feature replay and prompt tuning.
@@
Explicit Regularization. For explicit regularization, the central and challenging work is to estimate the importance of network parameters. After learning each task, the importance of individual parameters in the model is estimated to regularize the new class learning process by penalizing the change of important parameters. The general regularization loss is can be formalized as $\mathcal{L}_{\text{EReg}}^{t}=\sum_{i}^{P}\Omega_{i}(\theta_{i}^{t}-\theta_{i}^{t-1})^{2}$, where $\Omega_{i}$ is the estimated importance factor for parameter $\theta_{i}$, and $P$ is the number of parameters in the network. The first work to achieve this goal is elastic weight consolidation (EWC) [115], which uses the diagonal of the Fisher information matrix to compute the importance $\Omega_{i}$. Alternatively, parameter importance can be estimated by quantifying the sensitivity of the predicted output function to changes in each parameter [117]. These methods can be interpreted as metaplasticity-inspired [122] from a neuroscience point of view [123]. Nonetheless, designing a reasonable importance metric within a neural network remains challenging. Another line of works [124, 125, 126] avoid interfering with previously learned knowledge from a parameter space perspective by projecting and updating the parameters in the null space of the previous tasks. Those methods assume that the capacity of a neural network is high enough to learn new tasks effectively, which would not always hold in practice, especially for incremental learning with long sequences.
Implicit Regularization. Rather than directly constraining the parameters of the network, implicit regularization approaches focus on keeping the input-output behavior of the network based on current training data. The general regularization loss can be formalized as $\mathcal{L}_{\text{IReg}}^{t}=\mathbb{E}_{(\bm{x},y)\backsim\mathcal{D}^{t}}[D(f^{t-1}(\bm{x}),f^{t}(\bm{x}))]$, where $D(\cdot)$ denotes the distance metric such as Euclidean distance, Cosine distance, or Kullback-Leibler divergence. To realize this idea, Li et al. proposed learning without forgetting (LwF) which firstly incorporates the knowledge distillation (KD) [127] technique into incremental learning to consolidate old knowledge by distilling the outputs of the previous model to that of the updated model. Noteworthy, although KD [127] was originally developed for model compression, it has been widely used in CIL and served as a basic building block in many methods. Later, several works followed the teacher-student distillation framework to address the catastrophic forgetting in CIL by modifying the KD technique. For example, some approaches distill important features in the deep feature space [128, 129]. Specifically, important features can be identified via attention mechanisms [129, 130] or feature projection [128]. By distilling knowledge from important features, the model not only preserves the crucial information of old classes but also provides flexibility for adapting to new classes. Other works explore methods to distillate knowledge from multi-models [131, 132] or auxiliary data [131, 133, 134].
5.2 Exemplar Replay based CIL
Rebuffi et al. [135] firstly investigated exemplar replay strategy, and proposed iCaRL that learns both new data and stored old data at each incremental stage. Due to the strong performance, various exemplar replay methods have been developed, mainly focusing on four different aspects: classifier calibration, distilled knowledge, model structure and generative replay.
@@ 5.2 Exemplar Replay based CIL
Imbalance Calibration. Due to the memory limitation, the number of saved samples is usually much fewer compared with that of new classes. Therefore, there exists the class imbalance problem [136] between new and old classes. Visually, the norms of the weight vectors of old classes are much smaller than those of new classes [137]. To calibrate the bias, some approaches aim to learn a balanced classifier during training, while other approaches use post-hoc methods to calibrate the classifier. For training-time imbalance calibration, a classical way is under-sampling [138, 139], which reduces the number of new data to construct a balanced training set. However, this is undesirable because many samples of new classes remain unused. A better way is using the balanced training set to fine-tune the model with a small learning rate [140, 133, 141]. Recent work also shows that the memory management policy for imbalance calibration can also be learned via reinforcement learning. Besides focusing on data, other methods explore to modify the classifier directly such as cosine normalization [142] and separated softmax [143, 134]. For post-hoc imbalance calibration, iCaRL [135] introduces the nearest class mean classifier [59] to replace the original classifier, which effectively mitigates the class imbalance problem. A more direct way is to calibrate the classifier via bias correction layer [144] or output scaling [137, 145, 146]. For example, scaling the output logits based on logit norm [137] or rectifying softmax probabilities [146] for old classes to ensure a more balanced treatment of old and new classes during inference.
Distilled Knowledge. Knowledge distillation (KD) is a widely used technique to avoid catastrophic forgetting in exemplar replay methods, and it is often applied to both the saved old data and new data. A natural question is: What kind of knowledge should be distilled to maintain the performance on old classes? Indeed, many approaches have been developed from the perspective of distilled knowledge: some focus on unstructured KD, while more recent works focus on structured KD. For unstructured KD, early methods [135, 140, 144] adopt a straightforward and naive way that distills the soft probability distribution, while others [141] also prevent changes in the final or intermediate feature representations to maintain spatial information. To better preserve the previous knowledge, structural KD techniques penalize structural differences in relation of data examples. For instance, the neighborhood relations of samples [147] can be distilled to preserve the topology of the feature space during incremental learning. An alternative way is to penalize the change of ranking order of similarities [148] or instance-wise similarity [149]. Besides explicitly penalizing the change of neighborhood relation, GeoDL [150] models and punishes the gradual changes between two different tasks with the geodesic flow, while DDE [151] implicitly constrains the neighborhood relation by predicting a sample with its neighborhood.
Model Architecture. In task incremental learning (TIL) setting, dynamically extending the network architecture has been demonstrated to be quite effective [152, 153]. Specifically, when learning a new task, the old parameters are frozen and new branches are allocated to learn new tasks. However, those methods are often impractical for CIL because the task identity is not available at inference time, and one can not know which group of parameters should be used for inference given an input. Recently, based on exemplar replay, several architecture-based methods have been proposed for CIL. Liu et al. [154] designed two types of residual blocks that have different levels of learnability: a stable block is to maintain the knowledge of old classes, and a plastic block is to learn new classes. A similar idea has been explored in [155], which designed a dual networks architecture that includes a slow learning module for generic representation learning and a fast learning module for task-specific representation learning. Indeed, a simpler and more effective way is dynamically expanding feature extractor to preserve old knowledge while learning new concepts [156]. Specifically, at each incremental stage, the old feature extractor is frozen and a new feature extractor is introduced for learning new classes. However, incrementally creating new models leads to an increase in the number of parameters, while KD can be an effective way to remove redundant parameters [157].
@@
Generative Exemplar Replay. Despite the success of real exemplar replay methods in alleviating catastrophic forgetting, they may bring issues such as violation of data privacy and memory limitation: storing a fraction of old raw data would be unsuitable for privacy-sensitive applications in the field of healthcare and security [119, 158], and suffer from memory limitation for long-step incremental learning and applications with limited computational budget. Moreover, directly storing some old training samples is less human-like from the biological perspective [159, 160, 124]. To address privacy and memory limitation concerns of real data replay, researchers have started to consider generative data replay. Early approaches [161, 162] leverage generative model to generate pseudo-samples for old classes. Particularly, the generative model is trained simultaneously with the classification model at each incremental stage. However, those approaches perform poorly for CIL and the generated pseudo-examples rely heavily on the quality of the generative model. Besides, it is computational and memory intensive to train generative models, which also suffer from catastrophic forgetting in the incremental learning process. Recently, rather than leveraging an additional complex generative model, some works [163, 164] explore to use of the classification model itself to generate pseudo-samples. In those works, a crucial challenge is to address the distribution shift between real and generated samples.
5.3 Feature Replay based CIL

### v3

@@ V-A Problem Formulation and Key Challenges
However, directly updating the model on new classes would cause the catastrophic forgetting problem [340, 341, 342, 343], i.e., the classification accuracy on old classes drops dramatically after learning new classes. By taking a closer look at catastrophic forgetting in CL, we find that there are mainly three issues, and each of them represents one aspect of forgetting, while the concurrence of them results in the performance degradation of old classes after learning new classes in CL, as demonstrated as follows and illustrated in Fig. 8. (1) Representation and classifier mismatch: Without accessing old training data, directly updating the model on new classes could dramatically change the learned representations and class weights for old classes, which leads to the mismatch between old class features and old class weights. Consequently, the input-output behavior of old classes would be forgotten. (2) Representation confusion: After updating the model on new classes, the feature extractor would no longer be suitable for old classes. As a result, the input from old and new classes would be confused and overlapped in the deep feature space, making it hard to distinguish them at inference time. (2) Classifier imbalance: When learning new classes, the class weights of old and new classes can not be jointly optimized because of the lack of old data. Therefore, the class weights of old and new classes are often imbalanced, and the decision boundary can easily shift toward novel classes. The above analysis would not only give a deeper understanding of catastrophic forgetting in CL but also provide a unified perspective to understand existing methods.
V-B Regularization Methods
@@
1) Explicit Weight Constraint. For explicit regularization, the central and challenging work is to estimate the importance of network parameters. After learning each task, the importance of individual parameters in the model is estimated to regularize the new class learning process by penalizing the change of important parameters. The general regularization loss can be formalized as:
$\mathcal{L}_{\text{E-reg},t}=\sum_{i}^{P}\Omega_{i}(\bm{\Theta}_{i}^{t}-\bm{\Theta}_{i}^{t-1})^{2},$ (15)
@@
where $\Omega_{i}$ is the estimated importance factor for parameter $\bm{\Theta}_{i}$, and $P$ is the number of parameters in the network. As can be seen, the central and challenging work is to estimate the importance of network parameters. The first work to achieve this goal is elastic weight consolidation (EWC) [111], which uses the diagonal of the Fisher information matrix to compute the importance $\Omega_{i}$. Alternatively, parameter importance can be estimated by quantifying the sensitivity of the predicted output function to changes in each parameter [113]. Kong et al., [344] theoretically shown that the forgetting has a relationship with the maximum eigenvalue of the Hessian matrix, and proposed to explore the maximum eigenvalue to overcome catastrophic forgetting. Following those methods, various weight constraint approaches have been developed, such as RWalk [114], VCL [115], UCL [116], KCL [117] and AGS-CL [118]. These methods can be interpreted as metaplasticity-inspired [345] from a neuroscience point of view [159]. Nonetheless, designing a reasonable importance metric within a neural network remains challenging. Another line of works such as OWM [119], OGD [120], AOP [121], GPM [122], Adam-NSCL [123], NCL [124] and DSFGP [346] avoid interfering with previously learned knowledge from a parameter space perspective by projecting and updating the parameters in the orthogonal direction or the null space of the previous tasks. Those methods assume that the capacity of a neural network is high enough to learn new tasks effectively, which would not always hold in practice, especially for continual learning with long sequences.
Figure 9: The evolutionary tree of representative continual learning methods.
2) Implicit Knowledge Distillation. Rather than directly constraining the parameters of the network, implicit regularization approaches focus on keeping the input-output behavior of the network based on current training data. The general regularization loss can be formalized as:
$\mathcal{L}_{\text{I-reg},t}=\mathbb{E}_{(\bm{x},y)\backsim\mathcal{D}_{t}}[D(\text{out}^{t-1}(\bm{x}),\text{out}^{t}(\bm{x}))],$ (16)
where $D(\cdot)$ represents the distance metric such as Euclidean distance, $cos$ distance and Kullback-Leibler (KL) divergence. To realize this idea, Li et al. [347] proposed learning without forgetting (LwF), which firstly incorporates the knowledge distillation (KD) [126] technique into continual learning to consolidate old knowledge by distilling the outputs of the previous model to that of the updated model as follows:
$\mathcal{L}_{\text{LwF},t}=\mathbb{E}_{(\bm{x},y)\backsim\mathcal{D}_{t}}\left[-\sum_{i=1}^{|\mathcal{C}_{\text{old}}|}p_{i}^{t-1}(\bm{x})\text{log}~{}p_{i}^{t}(\bm{x})\right],$ (17)
where $|\mathcal{C}_{\text{old}}|=|\mathcal{C}_{1:t-1}|$ is the number of old classes and $p_{i}^{t-1}(\bm{x})$, $p_{i}^{t}(\bm{x})$ are the previous and current probabilities after temperature scaling. By optimizing $\mathcal{L}_{LwF,t}$, the KL divergence between the two probability distributions could be minimized. Noteworthy, although KD [126] was originally developed for model compression, it has been widely used in CL and served as a basic building block in many methods.
Figure 10: Illustration of regularization methods. (A) Explicit weight constraint. (B) Implicit knowledge distillation.
Later works followed the teacher-student distillation framework to address the catastrophic forgetting in CL by modifying the KD technique. Some approaches distill important features in the deep feature space [127, 128], where important features can be identified via attention mechanisms [128, 348] or feature projection [127]. For example, to evaluate the importance of features, encoder-based lifelong learning (EBLL) [127] projects the feature to a subspace with lower dimensions based on autoencoders. Then the Euclidean distance is used to distill those important features. Learning without memorization (LwM) [128] identifies the important features based on attention technique [348], and then penalizes the changes in classifiers’ attention maps to retain important information of the old classes when learning new classes. By distilling knowledge from important features, the model not only preserves the crucial information of old classes but also gives more space to adjust to the new classes. Other works explore methods to distillate knowledge from multi-models [129, 131] or auxiliary data [129, 349, 132]. For instance, deep model consolidation (DMC) [129] distills the knowledge from both the old model and the new model via a novel double distillation training objective. Instead of sequentially distilling knowledge only from the penultimate model, M2KD [130] directly leverages all previous models for KD. Alternatively, MUC [131] distills knowledge from an ensemble of auxiliary classifiers, to estimate more efficient and effective regularization constraints. Besides, some studies found that distillation on new classes can not retain old knowledge effectively, and the model could easily be biased towards new classes. To address this problem, Lee et al., [349] leverages a large stream of unlabeled auxiliary data for KD. CCIL et al. [132] adopts cutout [350] to synthesize novel data for KD, and demonstrated its effectiveness on calibrating the model’s predictions. However, as noticed by previous works [351, 352, 159], those regularization methods often failed dramatically for class continual learning scenario. Therefore, KD technique is often combined with data replay [133, 134, 135, 136, 137, 138], which will be detailed below.
V-C Data Replay Methods
A basic assumption in CL is the inaccessibility of old training data. However, data replay methods relax this constraint by saving and replaying some training samples of old classes,i.e., $\mathcal{M}$. Then the model is jointly trained on $\mathcal{M}\bigcup\mathcal{D}_{\text{t}}$. Intuitively, with replayed data, the previous knowledge would be consolidated, and the three problems (demonstrated in Fig. 8) in CL would be alleviated: First, the representation and classifier of old classes would be better matched because old samples are relearned in an end to end manner when learning new classes. Second, the representation confusion and class weights imbalance problems between old and new classes would be alleviated because data replay allows the discrimination between old and new classes. Rebuffi et al. [141] firstly investigated the data replay strategy for CL. Due to the strong performance, various data replay methods have been developed, mainly focusing on imbalance calibration and generative replay.
@@ V-C Data Replay Methods
1) Imbalance Calibration. Due to the memory limitation, the number of saved samples is usually much lower compared with that of new classes. Therefore, there exists the class imbalance problem [353] between new and old classes. Visually, the norms of the weight vectors of old classes are much smaller than those of new classes [143]. To calibrate the bias, some approaches aim to learn a balanced classifier during training, while other approaches use post-hoc methods to calibrate the classifier. A simple and classical way is under-sampling [354, 140]. For example, end-to-end incremental learning (EEIL) [139] includes additional balanced fine-tuning at the end of each incremental stage. Specifically, a balanced dataset that contains the same number of samples per class is constructed by reducing the number of samples from the new classes. Then, the balanced dataset is used to fine-tune the model with a small learning rate. Lee et al., [349] also adopt a balanced fine-tuning strategy by scaling the gradient based on the number of samples in each class. GDumb [140] greedily stores samples in memory and trains a model from scratch using samples only in the memory at inference time. These methods could alleviate the forgetting problem in CL, but lead to under-fitting of new classes since a lot of samples of new classes are not learned.
Other training-time imbalance calibration methods explore modifying the classifier directly, such as cosine normalization [133] and separated softmax [146, 132]. A representative method is UCIR [133], which adopts cosine normalization and margin ranking loss to learn a balanced classifier. On the one hand, UCIR normalizes both the input and class weights in the softmax options when computing the predicted probability of a sample:
$p(c|\bm{x})=\frac{\text{exp}{(\eta\langle\bar{f}_{\bm{\theta}}(\bm{x}),\bar{\bm{\varphi}}_{c}}\rangle)}{\sum_{j}\text{exp}{(\eta\langle\bar{f}_{\bm{\theta}}(\bm{x}),\bar{\bm{\varphi}}_{j}}\rangle)},$ (18)
[…]
$\mathcal{L}_{mr,t}(\bm{x})=\sum_{k=1}^{K}\text{max}~{}(m-\langle\bar{f}_{\bm{\theta}}(\bm{x}),\bar{\bm{\varphi}}_{y}\rangle+\langle\bar{f}_{\bm{\theta}}(\bm{x}),\bar{\bm{\varphi}}_{k}\rangle),$ (19)
where $m$ is the margin, $\bar{\bm{\varphi}}_{y}$ and $\bar{\bm{\varphi}}_{k}$ refer to the ground truth and the top-K new class embeddings, respectively. Those two strategies can effectively eliminate the classifier bias in CL. Ahn et al. [146] systematically analyzed the cause of classifier bias, and proposed a method that integrates separated softmax and task-wise knowledge distillation to resolve the bias. Those two strategies alleviate the bias accumulation in the CL process. Recently, Liu et al. [355] proposed a reinforcement learning based strategy to learn both the task-level and class-level memory allocation in different incremental phases to address the data imbalance problem.
Figure 11: Illustration of data replay methods. (A) Real data replay. (B) Generative data replay.
TABLE VIII: Summarization of imbalance calibration strategies.
Methods Imbalance Calibration Settings Imbalance Calibration Strategy
Data Classifier Train Test
EEIL [139] Two-stage: construct a balanced dataset to fine-tune the model.
GDumb [140] Down-sampling: build a balanced dataset and train the model from scratch.
SS-IL [146] Decouple softmax operations and apply knowledge distillation.
RMM [355] Reinforcement learning to optimize class balance in memory.
UCIR [133] Feature and weight normalization with ranking loss.
iCaRL [141] Classify based on nearest class mean.
BiC [142] Calibrate prediction bias using a separate balanced validation set.
WA [143] Align magnitudes of classification weights between new and old classes.
IL2M [145] Post-hoc calibration of model’s probability distribution.
ScaIL [144] Reshape classifier using aggregate class statistics.
ItO [147] Introduce deviation compensation to maintain learned decision boundary.
For post-hoc imbalance calibration, Rebuffi et al. [141] notice the classifier imbalance problem and introduce the nearest class mean (NCM) classifier at inference time. Technically, the NCM is computed by averaging the feature representations of each class:
$\bm{\varphi}_{c}\leftarrow\frac{1}{n_{c}}\sum_{i}\mathbb{I}(y_{i}=c)f_{\bm{\theta}}(\bm{x}_{i}),~{}~{}\forall c\in\mathcal{C},$ (20)
where $\mathcal{C}$ denotes all the encountered classes, $n_{c}$ denotes the number of the samples in class $c$. Particularly, for old classes, only the preserved data are used to compute the class mean. Compared with the original linear classifier, the NCM classifier is less sensitive to the class imbalance problem [49]. A more direct way is to calibrate the classifier via bias correction layer [142] or output scaling [143, 144, 145]. For instance, Bias Correction (BiC) [142] calibrates the output logits of the model by learning a linear bias correction layer:
$o^{\prime}_{c}\leftarrow\alpha o_{c}+\beta,~{}~{}\forall c\in\mathcal{C}_{\text{new}},$ (21)
where $o_{c}$ is the output logits for the new classes $c\in\mathcal{C}_{\text{new}}$. The bias calibration parameters i.e $\alpha$ and $\beta$ are estimated based on a balanced validation set and shared by all new classes. Weight Alignment (WA) [143] scales the output logits as follows:
$o^{\prime}_{c}\leftarrow\gamma o_{c},~{}~{}\forall c\in\mathcal{C}_{\text{new}},$ (22)
where
$\gamma=\frac{\text{Mean}((\|\bar{\bm{\varphi}}_{1}\|,...,\|\bar{\bm{\varphi}}_{|\mathcal{C}_{\text{old}}|}\|))}{\text{Mean}((\|\bar{\bm{\varphi}}_{|\mathcal{C}_{\text{old}}|+1}\|,...,\|\bar{\bm{\varphi}}_{|\mathcal{C}_{\text{old}}|+|\mathcal{C}_{\text{new}}|}\|))},$ (23)
in which Mean($\cdot$) returns the mean value of elements in the norm of classifier weights. Since scaling output logits is equivalent to scaling the norm of the weights, the method can be viewed as a strategy to align the average norm of the weight vectors for old and new classes. Similarly, ScaIL [144] also scales the classifier weights by leveraging statistical information of the model outputs. Instead of scaling the classifier weights or output logits, Belouadah et al. [145] rectified the output softmax probability by utilizing class statistics, i.e., the prediction score of the classifier. Concretely, if the prediction of an input $\bm{x}$ belongs to new classes, the softmax probability on any old class $c\in\mathcal{C}_{\text{old}}$ would be modified as follows:
$p^{\prime}(c|\bm{x})=p(c|\bm{x})\cdot\frac{\mu^{P}(c)}{\mu^{N}(c)}\cdot\frac{\mu^{N}(\mathcal{C}_{\text{new}})}{\mu^{P}(\mathcal{C}_{\text{old}})},\quad\forall c\in\mathcal{C}_{\text{old}},$ (24)
where $P$ denotes the state where the old class $c$ was learned initially; $N$ represents the current incremental state. $\mu^{P}(c)$ represents the mean prediction scores of all training samples of class $c$ at state P, and $\mu^{N}(c)$ represents the mean prediction scores of preserved samples of class $c$ at current state N. $\mu^{P}(\mathcal{C}_{\text{old}})$ and $\mu^{N}(\mathcal{C}_{\text{new}})$ share the similar meanings for all old classes $\mathcal{C}_{\text{ole}}$ and new classes $\mathcal{C}_{\text{new}}$, respectively. Table VIII provides a summary and comparison of different imbalance strategies in continual learning. Recently, Lai et al. [356] proposed a novel Pareto continual learning that reformulates the stability-plasticity trade-off in data replay based continual learning as a multi-objective optimization problem. Xu et al. [357] revealed the ineffectiveness of existing popular data replay methods and called for a more flexible, and practical CL setting where the storage buffer can be dynamically adjusted according to acceptable cost.
@@
2) Generative Data Replay. Despite the success of real data replay in alleviating catastrophic forgetting, there are issues such as violation of data privacy and memory limitation: storing a fraction of old raw data would be unsuitable for privacy-sensitive applications in the field of healthcare and security [358, 359], and suffer from memory limitation for long-step continual learning and applications with limited computational budget. Moreover, directly storing some old training samples is less human-like from the biological perspective [360, 361, 119]. To address privacy and memory limitation concerns of real data replay, researchers have started to consider generative data replay. Early approaches [148, 149, 152] leverage the generative model (e.g. missing, GAN [289, 362] and autoencoder [363]) to generate pseudo-samples for old classes. Recent investigations explore newly developed diffusion models [364] for replay, [155, 156, 154], enabling bidirectional instruction between generator and classifier to mitigate forgetting. Particularly, the generative model is trained simultaneously with the classification model at each incremental stage. However, those approaches perform poorly for CL, and the generated pseudo-examples rely heavily on the quality of the generative model. Moreover, it is computational and memory-intensive to train generative models, and they also suffer from catastrophic forgetting during continual learning. Rather than leveraging additional complex generative models, some recent works [365, 151, 366] explore to use of the classification model itself to generate pseudo-samples. Yin et al. propose the DeepInversion [365] to synthesize and replay pseudo-samples of previous classes when learning new classes. However, as noticed by [151, 366], the feature distribution of synthetic old samples is severely mismatched with that of real old samples, which introduces bias and misleads the decision boundary between old and new classes. Always Be Dreaming (ABD) [151] mitigates this issue by using separated softmax and balanced finetuning. Gao et al., [366] further improve ABD by leveraging relational knowledge distillation [367], which reduces forgetting while improving plasticity.
V-D Feature Replay Methods